# Bi-ConvLSTM: An Ultra-Lightweight Efficient Model for Human Activity Recognition on Resource Constrained Devices

### v1

@@ VIII Conclusion
Bi-ConvLSTM is an ultra-lightweight HAR architecture achieving 2.9 parameter reduction as compared to TinierHAR (11.4K vs. 34K parameters on average) while maintaining competitive accuracy within the ultra-lightweight regime. Key design choices include aggressive 4 temporal pooling, single layer bidirectional LSTM, and last timestep aggregation, enabling a 23.0 KB INT8 average deployment footprint with only 0.21% quantization degradation.
Our systematic evaluation reveals both strengths and limitations. Bi-ConvLSTM excels on smartphone IMU datasets (within 3% of larger baselines) and binary medical classification, but shows sensitivity to severe class imbalance on datasets like PAMAP2. Ablation studies demonstrate that architectural contributions are task dependent: bidirectionality benefits episodic event detection, but provides marginal or negative impact on periodic activities.
These findings suggest that the ultra-lightweight regime (15K parameters) is viable for single sensor wearable applications, but requires careful task architecture matching. Future work will investigate Neural Architecture Search for sensor modality specific configurations, and knowledge distillation to recover accuracy gaps while preserving the compact footprint.

### v3

@@ VIII Conclusion
Bi-ConvLSTM is an ultra-lightweight HAR architecture achieving 2.9 parameter reduction as compared to TinierHAR (11.4K vs. 34K parameters on average) while maintaining competitive accuracy within the ultra-lightweight regime. Key design choices, namely aggressive 4 temporal pooling, single-layer bidirectional LSTM, and last timestep aggregation, yield a compact model that deploys successfully on both the Raspberry Pi Pico 2 and ESP32 across all eight benchmarks under INT8 quantization.
On-device validation confirms that Bi-ConvLSTM is the only architecture in this study achieving full or near full dataset coverage on both target MCUs under both precision paths. Under INT8 quantization, it achieves 8/8 coverage on both platforms with 72.8 ms average latency and 85.7% PyTorch parity on Pico 2, and 168.5 ms and 97.9% parity on ESP32. Under FP32 full precision deployment, it achieves 8/8 on Pico 2 and 7/8 on ESP32 with 100.0% parity on all successful configurations, confirming that all INT8 fidelity degradation is a quantization artifact rather than an architectural limitation. By contrast, DeepConvLSTM (136K parameters) fails to deploy on either platform under both precision paths, while TinyHAR and TinierHAR achieve only partial coverage with considerable parity degradation under INT8 quantization.
Ablation studies reveal that architectural contributions are task dependent: bidirectionality benefits episodic event detection but provides marginal or negative impact on periodic activities. The cross platform INT8 parity discrepancy (85.7% vs. 97.9%) exposes an underreported concern in TinyML research, namely that desktop INT8 evaluation does not uniformly predict on-device fidelity across MCU backends. The FP32 results eliminate this variation entirely, establishing quantization as the dominant fidelity bottleneck.
These findings suggest that the ultra-lightweight regime (under 15K parameters) is viable for single sensor wearable applications when accompanied by rigorous on-device validation across both quantized and full-precision paths. Future work will investigate Neural Architecture Search for sensor-modality-specific configurations, knowledge distillation to recover INT8 parity gaps, and systematic characterization of backend dependent quantization fidelity across additional MCU targets. Code with PyTorch models, and quantized models: https://github.com/WhiteMetagross/MicroBiConvLSTM